Science Report – November 2nd

Crew 214 Science Report 2Nov2019
SOL: 6
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: Preliminary viewing, through the binocular microscope, of the particles recovered from yesterday’s EVA clearly showed the morphology of the particles and that some were spherical, however positive identification of micrometeorites could not be confirmed.
The fourth of the scientific EVAs was scheduled for the afternoon work period. This involved driving to site #2 at Tank Wash and continuing sample collection from the 1m x 1m plots. During the 2 hour EVA, 10 samples were collected and all appear to contain magnetic material.
On the return to the hab, the Cactus Road #1 site was revisited to allow another crew member an opportunity to operate the heliograph. With line of sight radio communications established, the heliograph was successfully operated and video records collected.
Tabulation of the newly arrived food stock macro and micronutrients for the waste study was completed.
Closed Loop Food Waste Study: The day’s waste for Sol 6 is: 15 grams; consisting of 44 calories, 1g fat, >1g saturated fat, 16mg sodium, 7g carbohydrates, 1g fibre, >1g sugar, 6g protein. At the end of the first week of the rotation, over 613 grams (2791 calories) of food have been wasted at a rate of approximately 400 calories per day, or 10 grams per person per day. This waste was proportionally derived from carbohydrates (63%), protein (15%) and fats (22%), which was a slight shift from the macronutrient ratios of crew stock inventory from the start of the rotation (C58%, P12%, F31%).
Additional Activities: Tidying and cleaning chores around the hab.

Journalist Report – November 2nd

Crew 214 Sol 6 2 Nov 2019
by Sandy Dance
Following our water scare the other day, we decided to delay the pumping of water from the low tank to the high tank for as long as possible. Consequently, despite the CapCom advising that we pump the water up last night we decided to wait until this morning. Mistake? Yes and no. When we came to pump this morning, nothing happened. So we contact the engineering team. The problem is likely that the pipes froze overnight, hence the caution in the MDRS Handbook on page 24 advising that the pumping happen in the evening. So we wait until noon. Still no water. This time the engineering team find that the actual(?) problem was that the hot water system was plugged into the same circuit, overloading it. Now we have water! And the lesson: this hab is complex, and needs to be treated with respect.
Today the EVA expedition went out to continue the micrometeorite experiment, scanning the square metre plots for micrometeorites. They were able to complete another 10 plots, finding more fine magnetic dust in each.
This morning we inspected yesterday’s findings with a 40X optical microscope, but we are still unable to determine the origin of the particles. We may need to use an electron microscope back in Australia for this.

Sol Summary – November 2nd

Sol:6

Summary Title: Boomer 1 & 2 go on EVA
Author’s name: Guy Murphy
Mission Status: Active.
Sol Activity Summary: The morning was spent attending to hab maintenance, cleaning tasks. During the afternoon, two crew members (Boomer 1 and 2) undertook an EVA to continue collecting micrometeorite samples. 3 meals were prepared today, including toasted sandwiches with Spam & Polish pickels.
Look Ahead Plan: Sunday 3 November (Sol 7) is a rostered day of rest for the crew, with no EVA’s scheduled. We will resume our normal work schedule on Monday 4 November.
Anomalies in work: None.
Weather: Clear skies, no precipitation. Below freezing overnight.
Crew Physical Status: All in good health.
EVA:Afternoon EVA completed successfully, no problems encountered.
Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.
Support Requested: No further support requested. The crew is enjoying the new dietary items that were provided to us in yesterday’s delivery.

Sol Summary – November 1st

Sol:5

Summary Title: About the Science
Author’s name: Guy Murphy
Mission Status: Active.
Sol Activity Summary: Deliveries of both hardware and food supplies arrived today, many thanks to Capcom for arranging these. Today was about the science, which we hope our selection of photographs illustrate. Key milestones were reached in two research projects, with collection of our first magnetic samples from the field as part of the micro-meteorite study and the cataloguing of the remaining food inventory by Dianne McGrath. Two cooked meals prepared today, with a cold breakfast. Routine internal hab cleaning was undertaken during the afternoon.
Look Ahead Plan: We hope to complete a large proportion of our planned micro-meteorite collection tomorrow.
Anomalies in work: None.
Weather: Clear skies all day, no wind, ideal for field work. Crew Physical Status: The crew continue to be in excellent health EVA: Successfully undertaken, see EVA report.
Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.
Support Requested: Some milk would be appreciated if there are any further incoming deliveries. We do have alternative sources of calcium (including 2 cans dried cheese), so the lack of milk will not be detrimental to bone density for the foreseeable future.

Science Report – November 1st

Crew 214 Science Report 1 November 2019
SOL: 5
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The third of the scientific EVAs was scheduled for the afternoon work period. This involved driving to site #2 at Tank Wash and continuing setting out the 100 square metre (100 1m x 1m plots). During the 2 hour EVA, setting out of the grid was completed and the first samples collected. This involved hovering the heavy duty magnet just above the ground surface and systematically ‘sweeping’ it across the 1 square metre of the plot. Two plots were sampled and very fine magnetic particles were recovered from both plots!
On the return to the hab, the Cactus Road #1 site was revisited to allow another crew member an opportunity to operate the heliograph. With line of sight radio communications established, the heliograph was successfully operated and video records collected.
Tabulation of the baseline food stock macro and micronutrients for the waste study is being updated with the arrival of further supplies. Closed Loop Food Waste Study: The day’s waste for Sol 5 is: 108 grams; consisting of 383 calories, 6g fat, 3g saturated fat, 176mg sodium, 53g carbohydrates, 4g fibre, 3g sugar, 24g protein. The bread production errors were turned into tasty crackers/cookies (#nowaste).
Additional Activities: Nil.

Journalist Report – November 1st

1 Nov 2019 Sol 5 Journalist Report
by Sandy Dance
Last night was a very warm night, largely due to the filter in the heater having been changed yesterday, allowing it run far more efficiently. Today is a very normal day, we have settlled into a routine, the various science projects now proceeding apace.
For instance: at breakfast, lunch and dinner, Dianne our nutrition scientist is beavering away studying the total nutrient loss from food waste. Consequently she carefully weighs our dirty dishes and cooking pots, and again after they are clean and finds the difference. Its a remarkably painstaking task, and also requires a lot of data entry. Its really interesting how much food is wasted, in preparation, plate waste and spoilage. So far total waste is 31 gm per person per day. Total calories wasted all up is 400 calories per day for the group! This would be a significant loss on Mars.
The other major science project, the micrometeorite study, had a big step forward this afternoon. Andrew and Dianne went on EVA and finished the 10 metre grid installation, and began the magnetic sweep. They managed the first two squares (2 m^2) and got a lot of magnetic particles. We are cock a hoop with this development, although since the particles are so tiny, its hard to tell whether they are volcanic magnetite or micrometeorites. It requires a microscope to make the distinction: the meteorites should be spherical, the magnetite shards. We await the chance to have a look.
So after a relaxing meal of chicken, kidney beans, rice and vegetables, we settle back for the evening.

Operations Report – November 1st

Crew 214 Operations Report 01-11-2019

SOL: 5

Name of person filing report: David Mateus

Non-nominal systems: Opportunity, suit 10

Notes on non-nominal systems: Oppy brakes do not work, Suit 10 is not charging

Generator: run

Hours run: 14

From what time last night: 17:30

To what time this morning: 7:30

List any additional daytime hours when the generator was run: none

Solar— SOC % (Before generator is run at night): 71%

Diesel Reading –71 %

Station Propane Reading – 85 %

Ethanol Free Gasoline – N/A

Water (loft tank): 17 gallons

Water Meter: 1455145 units

Water (static tank) – 301 gallons

Static to Loft Pump used – No

Water in GreenHab – 150 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: It is not in the hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: tes

Hours: (before EVA): 106.6 h

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging):50%

Currently charging: Yes

Opportunity rover used: No

Hours: 65.5 h

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 112.5

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging):100%

Currently charging: Yes

Notes on rovers: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) No

Reason for use: None

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? Used by assistant director to go to town

CrewCar used and why, where? No used

General notes and comments: None

Summary of internet: functional

Summary of suits and radios: used for EVA, Suit 10 is not charging and we need batteries for 6 radios

Summary of Hab operations: All systems are nominal

Summary of GreenHab operations: No used

Summary of ScienceDome operations: All systems are nominal

Summary of RAM operations: No used

Summary of any observatory issues: No used

Summary of health and safety issues: All the crew are healthy

Questions, concerns and requests to Mission Support: We need new batteries for six radios.

EVA Report – November 1st

Crew 214 EVA Report 01-11-2019
EVA # 5
Author: Andrew Wheeler
Purpose of EVA: Geological Sampling
Start time: 1400
End time: 1600
Narrative: The third EVA for science saw us returning to the Tank Wash location #2 at 518477E 4253767N where we continued setting out the 100 square metre (100 1m x 1m plots) sampling grid. Having completed that task, we took our first samples looking for micrometeorites in the evaluation of the micrometeorite flux. 2 plots were sampled and very fine magnetic particles were recovered from both.
On the return to the hab, a revisit to the line of sight from the Cactus Road #1 location at 519570E 4251974N allowed another crew member exposure to the heliograph. With radio communications established and the hab personnel alerted, the heliograph operated as expected and was observed and recorded at the hab.
Destination: Drove from the hab along Cow Dung Road to Tank Wash and then returned to Galileo Road and the junction it has with Cactus Road. Coordinates (use UTM NAD27 CONUS): 518477E 4253767N
Participants: Andrew Wheeler, Dianne McGrath
Road(s) and routes per MDRS Map: Cow Dung Road to Tank Wash